OpenStudy (anonymous):

What is the value of f(8) for the function f(x) = 3x – 2?

OpenStudy (ash2326):

@Hoa Please don't just give answer. Explain how you got 22

OpenStudy (anonymous):

yes i like to learn thank you

OpenStudy (ryan123345):

plug in 8 for the x \[f(8)=3(8)-2\]\[f(8)=24-2\]\[f(8)=22\]

OpenStudy (ash2326):

@InsanelyChaotic do you understand?

OpenStudy (anonymous):

oh my god yes...thats a lot more simple than i had though. thank you.